OpenStudy (anonymous):

does anyonee know how to simplify radical exponents, with variables (Algebra)

OpenStudy (anonymous):

google it

OpenStudy (anonymous):

\[\sqrt{5}\times \sqrt{30}\]

OpenStudy (anonymous):

Example question?

OpenStudy (saifoo.khan):

5 root6

OpenStudy (anonymous):

how did you get 5root6

OpenStudy (anonymous):

wait is it 5root6 or root35?!

OpenStudy (saifoo.khan):

y calc gave me that ans, lol.

OpenStudy (saifoo.khan):

my*

OpenStudy (saifoo.khan):

i knw how to work that out too.

OpenStudy (anonymous):

@saifoo.khan : THANKS:D

OpenStudy (anonymous):

\[\sqrt{30} *\sqrt{5}=\sqrt{30* 5}\]

OpenStudy (anonymous):

@imranmeah91: i dont think that right! ive never seen it done that way! lol

OpenStudy (saifoo.khan):

root150. root150 can be written as root 25*6 therefore 5root6

OpenStudy (anonymous):

OOOH!

OpenStudy (saifoo.khan):

barbie, he was also right, lol.

OpenStudy (anonymous):

That's only way to figure it out

OpenStudy (anonymous):

hahaha oh wellthanks you guysss!

OpenStudy (saifoo.khan):

lol.